a fruit drink concentrate is 15 % water. How many liters of pure water should be added to 12 liters of the concentrate to get a mixture that is 50 % water?
Answer by checkley77(12844)

.15*12+W=.5(12+W)
1.8+W=6+.5W
W-.5W=6-1.8
.5W=4.2
W=4.2/.5
W=8.4 LITERS OF PURE WATER IS USED.
PROOF:
.15*12+8.4=.5(12+8.4)
1.8+8.4=.5*20.4
10.2=10.2